# Overlap-aware End-to-End Supervised Hierarchical Graph Clustering for Speaker Diarization

### v1

@@ III-E2 Clustering
III-E3 Feature aggregation
The cluster nodes at level $m$ are aggregated to form node features of the next level. To obtain node representations for next level $\boldsymbol{\mathbb{H}}_{t}^{m+1}$, the clusters $\bm{\mathcal{C}}_{t}^{m}$ and the features $\boldsymbol{\mathbb{H}}_{t}^{m}$ are used to compute identity feature ${\tilde{\bm{h}}_{i}}^{(m+1)}$ and average feature ${\bar{\bar{\bm{h}}}_{i}}^{(m+1)}$ of each cluster $i$ as,
$\begin{split}{\tilde{\bm{h}}_{i}}^{(m+1)}&={\tilde{\bm{h}}^{(m)}}_{z_{i}};\quad\quad{\bar{\bar{\bm{h}}}_{i}}^{(m+1)}=\frac{1}{|{\mathcal{C}_{i}}^{m}|}\sum_{j\in{\mathcal{C}_{i}}^{m}}{\tilde{\bm{h}}_{j}}^{(m)}\\ {\bm{h}_{i}}^{(m+1)}&=[{\tilde{\bm{h}}_{i}}^{(m+1)};{\bar{\bar{\bm{h}}}_{i}}^{(m+1)}]\end{split}$ (3)
@@ III-E3 Feature aggregation
III-E4 Graph generation
A new graph $G_{t}^{m+1}$ is constructed for the next level using the node features $\boldsymbol{\mathbb{H}}_{t}^{(m+1)}$. The edges are formed using $\bm{S}_{t}^{m+1}$ which is computed using the identity features $\boldsymbol{\mathbb{H}}_{t}^{(m+1)}$.
@@ III-E4 Graph generation
Figure 2: Block schematic of the E-SHARC training. The ETDNN and GNN modules in blue blocks contain learnable parameters. The GNN module generates edge probabilities and weights used in loss computation.
III-F Model training
E-SHARC training involves learning the embedding extractor and the GNN module (SHARC). During training, multiple graphs are constructed using input features and similarity matrix at different levels of clustering for each of the audio recordings in the training set. The loss combines binary cross entropy loss and mean squared error loss. These losses across all graphs are then accumulated and back-propagated for each batch. The block diagram is shown in Figure 2.
III-F1 Dataset generation
For the training set, input graphs $G=\{G^{0}_{1},G^{0}_{2},...,G^{1}_{1},...,G^{Mr}_{R}\}$ are constructed at different clustering levels for each recording $r$. $M_{r}$ is the maximum number of levels created for $r$. $E=\{E_{1}^{0},..,E_{1}^{M_{1}},...,E_{R}^{0},..,E_{R}^{M_{R}}\}$ and $V=\{V_{1}^{0},..,V_{1}^{M_{1}},...,V_{R}^{0},..,V_{R}^{M_{R}}\}$ are the set of all possible edges and nodes, respectively. For each level, clustering is performed based on edge probabilities, allowing only the same speaker embeddings to belong to a single cluster.
III-F2 SHARC training
This step uses a pre-trained extended time delay neural network (ETDNN) [27] model for embedding extraction to extract x-vectors. These x-vectors are used to generate the training graphs. In the first step, the weights of the ETDNN module are frozen, and the GNN module weights are learned.
@@ III-F2 SHARC training
Loss function: The GNN module is trained using the following loss function.
@@
where $L_{conn}$ is the pairwise binary cross entropy loss based on edge probabilities across all the possible edges in $E$ given as:
$L_{conn}=-\frac{1}{|E|}\sum_{(v_{i},v_{j})\in E}l_{ij}$ (5)
@@
Here, $|E|$ represents the total number of edges. $L_{den}$ represents the neighborhood density average loss given by Equation 7. $L_{den}$ represents mean squared error (MSE) loss between ground truth node density $d_{i}$ and predicted density $\hat{d}_{i}$ ,
$L_{den}=\frac{1}{|V|}\sum_{i=1}^{|V|}||d_{i}-\hat{d}_{i}||_{2}^{2}$ (7)
where $|V|$ is the cardinality of $V$.
III-F3 Joint ETDNN and GNN training
In this step, the GNN module is first initialized using the SHARC trained model. The ETDNN model is also learned along with the GNN module.
ETDNN architecture: The $13$-layer ETDNN model follows the architecture described in [27, 29]. The input to the model is 40-D mel-spectrogram features extracted from each segment (1.5s) of the training recording followed by cepstral mean normalization. The model comprises four blocks containing TDNN and a fully connected layer of size 1024D. This is followed by two feed-forward layers containing $\{1024,2000\}$ units and a segment pooling layer of 4000D concatenating the mean and standard deviation of each segment from the previous layer. This is passed to an affine layer of 512D which is called the x-vector embedding. The model is initialized with a pre-trained model for speaker classification. These x-vectors are used to generate the training graphs as described in the Section III-F1. The loss is back propagated till the TDNN layers.
@@ IV Handling Overlapped Speech
We extend the E-SHARC model also to perform GNN-based overlap prediction called E-SHARC-Overlap. Our approach assumes that a maximum of two speakers are present in an overlapping region. We follow a two-pass approach to accurately identify the speakers in the overlapping region. The first pass comprises E-SHARC modeling on the clean segments (single speaker segments). The first speaker, referred to as the parent cluster, is selected for each segment based on the first-pass E-SHARC algorithm. The proposed overlap model is trained using overlapping and clean segments in the second pass. During training, the graph adjacency matrix is generated by connecting the nodes from one cluster to nodes from any other cluster except the parent cluster. This enables the overlap model to identify the second speaker in the overlapping regions. A small percentage of intra-cluster connections ($10\%$) are preserved randomly, enabling contrastive training. The k-nearest neighbors of each node are selected based on the final graph adjacency matrix. The training loss comprises BCE and MSE losses, similar to the original SHARC model (Equation 4).
IV-A Initialization
The SHARC-Overlap and E-SHARC-Overlap models are initialized with the pre-trained SHARC and E-SHARC models, respectively.
IV-B Inference
@@
IV-B1 First speaker assignment
First-pass clustering based on E-SHARC is performed to assign an initial parent cluster and to decide the number of speakers present in a recording, as described in Section III-E. The input x-vectors are the average of the pre-trained x-vectors and the E-SHARC x-vectors.
IV-B2 Second/Overlap speaker assignment
@@ V-A Datasets
  • The AMI dataset: The official speech recognition partition of the AMI dataset  [30] comprises training, development (dev), and evaluation (eval) sets consisting of $136$, $18$, and $16$ recordings sampled at $16$kHz, respectively. The single-distant microphone (SDM) condition of the AMI dataset is used for experiments. It contains single-channel microphone recordings. The AMI train set contains 75 hrs of labeled speech. The number of speakers and the duration of each recording ranges from 3-5 and $20$-$60$ mins, respectively.
  • The Voxconverse dataset: It is an audio-visual diarization dataset[31] consisting of multispeaker human speech recordings extracted from YouTube videos. It is divided into a development (dev) set and an evaluation (eval) set consisting of $216$ and $232$ recordings, respectively. The duration of a recording ranges from $22$-$1200$ s. The number of speakers per recording varies from $1$-$21$.
    As the Voxconverse dataset does not have a train set, the dataset used for training is simulated using Voxceleb 1 and 2 [32, 33] and Librispeech [34] following the recipe mentioned in [35]. We simulated $5000$ mixtures containing $2$-$5$ speakers with duration ranging from $150$-$440$ s. This generates $1000$ hrs of data with $6,023$ speakers.
  • The DISPLACE challenge dataset: The DISPLACE 2023 [36] dataset released as part of the challenge comprises single-channel far-field natural multilingual, multi-speaker conversational speech recordings. The development (dev) and evaluation phase 2 (eval) sets contain 27 recordings (15.5 hours) and 29 recordings (16 hours), respectively. The speakers in the dev and eval sets are mutually exclusive. The dataset contains conversations in Hindi, Kannada, Bengali, Malayalam, Telugu, Tamil, and Indian English. In each recording, the number of speakers varies from $3$-$5$, and the number of languages varies from $1$-$3$.
@@ V-B Baseline system
The baseline method is an x-vector clustering-based approach described in [37, 9]. First, the recording is divided into 1.5s segments with 0.75s shift. The 40D mel-spectrogram features are computed from each segment which is passed to the ETDNN model [27] to extract 512D x-vectors. The ETDNN model is trained on the VoxCeleb1 [32] and VoxCeleb2 [33] datasets, for the speaker identification task, to discriminate among the $7,146$ speakers. The whitening transform, length normalization, and recording level PCA (dimension=30) are applied to the x-vectors as pre-processing steps. These embeddings are used to compute the PLDA similarity score matrix and clustered to generate speaker labels for each segment. The PLDA model is trained using the x-vectors.
For the baseline system, we have used the two most popular clustering approaches - Agglomerative Hierarchical Clustering (AHC) [38] and Spectral Clustering (SC) [18]. The PLDA similarity scores are used directly to perform AHC. For SC, we convert the PLDA scores $s$ to $s^{\prime}\in[0,1]$ by applying sigmoid with temperature parameter $\tau=0.1$ (best value obtained from experimentation) as: $s^{\prime}=\frac{1}{1+exp(-s/\tau)}$. This helps to improve the performance without other preprocessing steps usually followed in previous implementations [6].
@@
TABLE II: Performance comparison based on cluster purity and coverage for Voxconverse dataset. SC stands for Spectral Clustering.
@@
V-C Implementation details
V-C1 Optimizer
The SHARC model is trained with a Stochastic Gradient Descent (SGD) optimizer with a learning rate $lr$=$0.01$ (for Voxconverse) and $lr$=$1e$-$3$ (for AMI) for $500$ epochs. Similarly, the E-SHARC is also trained with an SGD optimizer. In this case, the learning rate is $1e$-$6$ for the ETDNN model and $1e$-$3$ for the GNN module; the model is trained for 20 epochs. SHARC-Overlap is initialized with SHARC model weights and trained with $lr$=$1e$-$3$ for 100 epochs. E-SHARC-Overlap is initialized with E-SHARC model weights and trained for five epochs.
V-C2 Adjacency matrix
The graph adjacency matrix of each recording is obtained using a similarity matrix with k highest similarities in each row. At each level of the hierarchy, the x-vectors, also called the node identity features, are passed to the PLDA model to generate the similarity score matrix.
VI Results
The performance of the system is measured based on Diarization Error Rate (DER) [39]. It is the sum of false alarm, miss rate, and speaker confusion error. The implementation is based on dscore toolkit [^1] https://github.com/nryant/dscore. We have used ground truth speech activity decisions for evaluating AMI and Voxconverse datasets.
We also evaluate our trained model on the DISPLACE challenge dataset. As per challenge guidelines, we have used the DISPLACE challenge baseline SAD model. The model is based on TDNN architecture [36, 37] with speech and non-speech classification. The Voxconverse E-SHARC model is also used for DISPLACE evaluations as the DISPLACE dataset does not contain a training partition.
@@ VI Results
E-SHARC-Overlap + VBx 0.7 11.6 4.91 17.21 0.8 1.6 7.7 10.11 3.8 21.0 6.6 31.40
VI-A Comparison with the baseline systems

### v2

@@ III-D2 Clustering
III-D3 Feature aggregation
The cluster nodes at level $m$ are aggregated to form node features of the next level. To obtain node representations for next level $\boldsymbol{\mathbb{H}}_{t}^{(m+1)}$, the clusters $\boldsymbol{\mathbb{C}}_{t}^{(m)}$ and the features $\boldsymbol{\mathbb{H}}_{t}^{(m)}$ are used to compute identity feature ${\tilde{\bm{h}}_{i}}^{(m+1)}$ and average feature ${\bar{\bar{\bm{h}}}_{i}}^{(m+1)}$ of each cluster $i$ as,
$\begin{split}{\tilde{\bm{h}}_{i}}^{(m+1)}&={\tilde{\bm{h}}^{(m)}}_{z_{i}};\quad\quad{\bar{\bar{\bm{h}}}_{i}}^{(m+1)}=\frac{1}{|{\bm{\mathcal{C}}_{i}}^{(m)}|}\sum_{j\in{\bm{\mathcal{C}}_{i}}^{(m)}}{\tilde{\bm{h}}_{j}}^{(m)}\\ {\bm{h}_{i}}^{(m+1)}&=[{\tilde{\bm{h}}_{i}}^{(m+1)};{\bar{\bar{\bm{h}}}_{i}}^{(m+1)}]\end{split}$ (3)
@@ III-D3 Feature aggregation
III-D4 Graph generation
A new graph $\boldsymbol{\mathbb{G}}_{t}^{(m+1)}$ is constructed for the next level using the node features $\boldsymbol{\mathbb{H}}_{t}^{(m+1)}$. The edges are formed using $\bm{S}_{t}^{(m+1)}$ which is computed using the identity features $\boldsymbol{\mathbb{H}}_{t}^{(m+1)}$.
@@ III-D4 Graph generation
Figure 2: Block schematic of the E-SHARC training. The ETDNN and GNN modules in black blocks contain learnable parameters. The GNN module generates edge probabilities and weights used in loss computation.
III-E Model training
The stages involved in E-SHARC model training are,
  1. 1.
    Pre-training - Train the x-vector module using speaker classification loss. Train PLDA using the x-vector embeddings and the speaker labels.
  2. 2.
    Data preparation - Preparing the training data in the form of graph inputs and outputs for GNN module training using x-vectors as node vectors and edge connections defined using PLDA.
  3. 3.
    GNN module training - Train the GNN module using the SHARC training procedure.
  4. 4.
    E-SHARC training - Train the joint model of x-vector embedding network and the GNN using the graph clustering loss.
The stage 1 is same as our baseline system which is described in Section V-B and the stages 2-4 are described in detail below.
Data preparation - Multiple graphs are constructed using input features and similarity matrices at different levels of clustering for each of the audio recordings in the training set. The loss combines binary cross entropy loss and mean squared error loss. These losses across all graphs are then accumulated and back-propagated for each batch. The block diagram is shown in Figure 2.
For the training set, input graphs $\boldsymbol{\mathbb{G}}=\{\boldsymbol{\mathbb{G}}^{(0)}_{1},\boldsymbol{\mathbb{G}}^{(1)}_{1},...,\boldsymbol{\mathbb{G}}^{(1)}_{1},...,\boldsymbol{\mathbb{G}}^{(M_{r})}_{R}\}$ are constructed at different clustering levels for each recording $r$. $M_{r}$ is the maximum number of levels created for $r$. $\boldsymbol{\mathbb{E}}=\{\boldsymbol{\mathbb{E}}_{1}^{(0)},..,\boldsymbol{\mathbb{E}}_{1}^{(M_{1})},...,\boldsymbol{\mathbb{E}}_{R}^{(0)},..,\boldsymbol{\mathbb{E}}_{R}^{(M_{R})}\}$ and $\boldsymbol{\mathbb{V}}=\{\boldsymbol{\mathbb{V}}_{1}^{(0)},..,\boldsymbol{\mathbb{V}}_{1}^{(M_{1})},...,\boldsymbol{\mathbb{V}}_{R}^{(0)},..,\boldsymbol{\mathbb{V}}_{R}^{(M_{R})}\}$ are the set of all possible edges and nodes, respectively. For each level $m$, graph $\boldsymbol{\mathbb{G}}^{(m)}_{r}$ consists of nodes that belong to the same speaker cluster and satisfies the condition in Equation 2 with $\tau=0$. The node features are obtained using Equation 3.
GNN module training - This step uses a pre-trained extended time delay neural network (ETDNN) [16] model for embedding extraction to extract x-vectors. These x-vectors are used to generate the training graphs. In the first step, the weights of the ETDNN module are frozen, and the GNN module weights are learned.
@@ III-E Model training
The GNN module is trained using the following loss function.
@@
where $L_{conn}$ is the pairwise binary cross entropy loss based on edge probabilities across all the possible edges in $\boldsymbol{\mathbb{E}}$ given as (as described in [15]) :
$L_{conn}=-\frac{1}{|\boldsymbol{\mathbb{E}}|}\sum_{(v_{i},v_{j})\in\boldsymbol{\mathbb{E}}}l_{ij}$ (5)
@@
Here, $|\boldsymbol{\mathbb{E}}|$ represents the total number of edges. $L_{den}$ represents the neighborhood density average loss given by Equation 7. $L_{den}$ represents mean squared error (MSE) loss between ground truth node density $d_{i}$ and predicted density $\hat{d}_{i}$,
$L_{den}=\frac{1}{|\boldsymbol{\mathbb{V}}|}\sum_{i=1}^{|\boldsymbol{\mathbb{V}}|}||d_{i}-\hat{d}_{i}||_{2}^{2}$ (7)
where $|\boldsymbol{\mathbb{V}}|$ is the cardinality of $\boldsymbol{\mathbb{V}}$.
E-SHARC training - All the model parameters including the $13$-layer ETDNN model, and the GNN model are fine-tuned using the final loss (Eq. 4). Following the E-SHARC training, we do not retrain the PLDA model as this was not found to improve the results further. The continuous update of the PLDA model parameters along with the SHARC training may potentially improve the modeling. However, in this paper, we have not updated the PLDA model.
@@ IV Handling Overlapped Speech
We extend the E-SHARC model also to perform GNN-based overlap prediction called E-SHARC-Overlap. Our approach assumes that a maximum of two speakers are present in an overlapping region. We follow a two-pass approach to accurately identify the speakers in the overlapping region. During training, the first pass comprises E-SHARC modeling on the clean segments (single speaker segments). The first speaker, referred to as the parent cluster, is selected for each segment based on the first-pass E-SHARC algorithm. The proposed overlap model is trained using overlapping and clean segments in the second pass. The graph adjacency matrix is generated by connecting the nodes from one cluster to nodes from any other cluster except the parent cluster. This enables the overlap model to identify the second speaker in the overlapping regions. A small percentage of intra-cluster connections ($10\%$) are preserved randomly, enabling contrastive training. The k-nearest neighbors of each node are selected based on the final graph adjacency matrix. The training loss comprises BCE and MSE losses, similar to the original SHARC model (Equation 4).
Inference
@@
IV-1 First speaker assignment
First-pass clustering based on E-SHARC is performed to assign an initial parent cluster and to decide the number of speakers present in a recording, as described in Section III-D. The input x-vectors are the average of the pre-trained x-vectors and the E-SHARC x-vectors.
IV-2 Second/Overlap speaker assignment
@@ V-A Datasets
  • The AMI dataset: The official speech recognition partition of the AMI dataset  [35] comprises training, development (dev), and evaluation (eval) sets consisting of $136$, $18$, and $16$ recordings sampled at $16$kHz, respectively. The single-distant microphone (SDM) condition of the AMI dataset is used for experiments. It contains single-channel microphone recordings. The AMI train set contains 75 hrs of labeled speech. The number of speakers and the duration of each recording is in the range 3-5 and $20$-$60$ mins, respectively.
  • The Voxconverse dataset: It is an audio-visual diarization dataset[36] consisting of multispeaker speech recordings extracted from YouTube videos. It is divided into a development (dev) set and an evaluation (eval) set consisting of $216$ and $232$ recordings, respectively. The duration of a recording is in the range $22$-$1200$ s. The number of speakers per recording ranges from $1$ to $21$. The dataset used for training is simulated using Voxceleb 1 and Voxceleb 2 [37, 38] and Librispeech [39] following the recipe mentioned in [40]. We simulated $5000$ mixtures containing $2$-$5$ speakers with duration ranging from $150$ to $440$ s. This generates about $400$ hrs of data with $6,023$ speakers.
  • The DISPLACE challenge dataset: The DISPLACE 2023 [41] dataset released as part of the challenge comprises single-channel far-field natural multilingual, multi-speaker conversational speech recordings. The development (dev) and evaluation phase 2 (eval) sets contain 27 recordings (15.5 hours) and 29 recordings (16 hours), respectively. The speakers in the dev and eval sets are mutually exclusive. The dataset contains conversations in Hindi, Kannada, Bengali, Malayalam, Telugu, Tamil, and Indian English. In each recording, the number of speakers ranges from $3$ to $5$, and the number of languages ranges from $1$ to $3$.
@@ V-B Baseline system
The baseline method is an x-vector clustering-based approach described in [42, 11]. First, the recording is divided into 1.5s segments with 0.75s shift. The 40D mel-spectrogram features are computed from each segment which is passed to the ETDNN model [16] to extract 512D x-vectors. The ETDNN model is trained on the VoxCeleb1 [37] and VoxCeleb2 [38] datasets, for the speaker identification task, to discriminate among the $7,146$ speakers. The input to the model is 40-D mel-spectrogram features extracted from each segment (1.5s) of the training recording followed by cepstral mean normalization. The model comprises four blocks containing TDNN and a fully connected layer of size 1024D. This is followed by two feed-forward layers containing $\{1024,2000\}$ units and a segment pooling layer of 4000D concatenating the mean and standard deviation of each segment from the previous layer. This is passed to an affine layer of 512D which is called the x-vector embedding. The whitening transform, length normalization, and recording level PCA (dimension=30) are applied to the x-vectors as pre-processing steps. These embeddings are used to compute the PLDA similarity score matrix and clustered to generate speaker labels for each segment. The PLDA model is trained using the x-vectors.
For the baseline system, we have used the two most popular clustering approaches - Agglomerative Hierarchical Clustering (AHC) [23] and Spectral Clustering (SC) [22]. The PLDA similarity scores are used directly to perform AHC. For SC, we convert the PLDA scores $s$ to $s^{\prime}\in[0,1]$ by applying sigmoid with temperature parameter $\alpha=0.1$ (best value obtained from experimentation) as: $s^{\prime}=\frac{1}{1+exp(-s/\alpha)}$. This helps to improve the performance without other pre-processing steps usually followed in previous implementations [8].
@@
TABLE II: Performance comparison based on cluster purity and coverage for Voxconverse dataset.
@@
+ E-SHARC-Overlap + VBx 0.7 11.6 4.91 17.21 0.8 1.6 7.7 10.11 3.8 21.0 6.6 31.40
TABLE V: DER (%, with overlap $+$ without collar), DER* (%, without overlap $+$ with 0.25s collar) and DER** (%, with overlap $+$ with 0.25s collar) comparison with state-of-the-art on AMI SDM , Voxconverse, and DISPLACE 2023 (phase 2) Evaluation datasets. prop.: proposed
AMI SDM System VAD DER DER* DER**
x-vec+AHC+VBx [43] Oracle 27.4 12.6 -
SelfSup-PLDA-PIC +VBx [11] Oracle 23.8 5.5 -
GAE-based+ SC [44] Oracle - 5.5 -
GADEC-based [44] Oracle - 4.2 -
E-SHARC (prop.) Oracle 19.8 2.9 -
E-SHARC-Ovp +VBx (prop.) Oracle 17.2 2.6 -
Pyannote 2.1.1 [28] System 27.1 - -
Raj et al. [45] System 23.7 - -
Plaquet et al. [46] System 22.9 - -
ResNet x-vec.+VBx+OSD [20] System 34.6 - -
DiaPer EEND [20] System 37.5 - -
E-SHARC (prop.) System 28.2 10.7 -
+ E-SHARC-Ovp +VBx (prop.) System 25.7 10.5 -
Voxconverse System
GAE-based+ SC [44] Oracle - 8.0 -
GADEC-based [44] Oracle - 7.7 -
E-SHARC (prop.) Oracle 11.7 7.6 9.2
+ E-SHARC-Ovp +VBx (prop.) Oracle $\bm{10.1}$ 6.3 7.5
Pyannote 2.1.1 [28] System 11.2 - -
Plaquet et al. [46] System 10.4 - 5.8
Pyannote VoxSRC2023 [47] System - - 4.0
Krisp VoxSRC2023 [48] System - - 4.4
ResNet x-vec.+VBx+OSD [20] System - - 6.1
DiaPer EEND [20] System - - 18.2
E-SHARC (prop.) System 16.1 9.7 11.3
+ E-SHARC-Ovp +VBx (prop.) System 15.1 8.8 10.0
DISPLACE System
Baseline [41, 49] Baseline 32.2 - -
E-SHARC-Ovp +VBx (prop.) Baseline 31.4 - -
E-SHARC-Ovp +VBx (prop.) System 29.9 - -
Winning system [49] System 27.8 - -
V-C Implementation details
V-C1 Optimizer
The SHARC model [^1] https://github.com/prachiisc/SHARC.git is trained with a Stochastic Gradient Descent (SGD) optimizer with a learning rate $lr$=$0.01$ (for Voxconverse) and $lr$=$1e$-$3$ (for AMI) for $500$ epochs. Similarly, the E-SHARC is also trained with an SGD optimizer. In this case, the learning rate is $1e$-$6$ for the ETDNN module and $1e$-$3$ for the GNN module; the model is trained for 20 epochs. SHARC-Overlap is initialized with SHARC model weights and trained with $lr$=$1e$-$3$ for 100 epochs. E-SHARC-Overlap is initialized with E-SHARC model weights and trained for $5$ additional epochs.
V-C2 Adjacency matrix
The graph adjacency matrix of each recording is obtained using a similarity matrix with k highest similarities in each row. At each level of the hierarchy, the x-vectors, also called the node identity features, are passed to the PLDA model followed by sigmoid activation to generate the similarity score matrix.
VI Results
The performance of the system is measured based on Diarization Error Rate (DER) [50]. It is the sum of false alarm, miss rate, and speaker confusion error. The implementation is based on dscore toolkit [^2] https://github.com/nryant/dscore. We have used ground truth speech activity decisions for evaluating AMI and Voxconverse datasets. However, we have integrated voice activity detection (VAD) system called Pyannote VAD [28] in our final model to compare with the other state-of-the-art approaches.
We also evaluate our trained model on the DISPLACE challenge dataset. As per challenge guidelines, we have used the DISPLACE challenge baseline VAD model. The model is based on TDNN architecture [41, 42] with speech and non-speech classification. The Voxconverse E-SHARC model is also used for DISPLACE evaluations as the DISPLACE dataset does not contain a training partition.
VI-A Comparison with baseline systems